# GROOD: Gradient-Aware Out-of-Distribution Detection

### v2

@@ 1 Introduction
To address these challenges, we propose a novel post-hoc method called GRadient-aware Out-Of-Distribution detection (GROOD) that relies on feature and gradient spaces for improved OOD discrimination. Our approach is built upon key insights from recent advancements in understanding neural network behavior, particularly the concept of neural collapse (NC(Papyan et al., 2020). This property suggests that prototype-based methods, like GROOD, can be highly effective when in-distribution data exhibits well-defined cluster structure. We leverage these insights to develop a more effective way of distinguishing between ID and OOD samples in the feature space of pre-trained networks. The core idea behind GROOD is to examine how sensitive the distances between sample representations and class prototypes are to changes in an artificially introduced OOD prototype. By analyzing this sensitivity through gradients, we can effectively differentiate between ID and OOD samples without the need for additional training.
Our method is inspired by two key observations:
  1. 1.
@@
    The NC property (Papyan et al., 2020) suggests that at the end of neural network training, within-class variability tends to zero for sample representation in the feature space. This observation motivates the use of nearest class prototype (NCP) classification, which relies on the distance to class prototypes, defined as the means of the samples of each class in the feature space. To further enhance the discriminative power of this approach for out-of-distribution detection, we construct logits by incorporating distances to an additional OOD prototype, alongside the distances to the class prototypes.
  2. 2.
@@
    We also observe that OOD samples tend to exhibit a more dispersed distribution in the feature space compared to ID samples. Capitalizing on this characteristic, we introduce an artificial OOD prototype, strategically positioned to be distinct from the ID class prototypes. By then examining how sample representations respond to this OOD prototype, specifically by analyzing gradients of the NCP loss with respect to it, we can gain a more nuanced understanding of the differences between ID and OOD samples, leading to more effective discrimination in the gradient space.
@@
Our approach differs from traditional post-hoc methods by computing gradients with respect to an artificial OOD prototype rather than the network’s parameters. The magnitude of these gradients serves as a key indicator: for ID data, the OOD prototype has a relatively small influence on the confidence of the prediction in the feature space, reflecting stable classification. In contrast, for OOD data, the OOD prototype has a more substantial influence on the confidence of the prediction; meaning that a smaller shift in the OOD prototype’s representation is sufficient to cause a larger change in the classification confidence.
We conduct an extensive empirical study following the recent methodology introduced in the OpenOOD Benchmark (Zhang et al., 2023a), but we also evaluate our method on other recent architectures.
@@ 3.2 Problem Setting: Out-of-Distribution Detection
[…]
@@ 4 GROOD Methodology
Figure 1: Initially, we build ID class prototypes as the means of the activations of ID data along with an OOD prototype capturing OOD characteristics (§ 5). Subsequently, gradients of the softmax loss built upon the NCPs distance as logits are computed w.r.t. OOD prototype (§ 4.1). Finally, the OOD score is determined using nearest neighbor distance in the gradient space (§ 4.2).
@@
In this section, we introduce our proposed method GRadient-aware Out-Of-Distribution detection (GROOD), a novel framework for distinguishing between ID and OOD samples. To illustrate the core mechanism, we will initially assume the existence of a representative OOD prototype, the calculation of which will be detailed in § 5.1. Notably, our oracle experiments (§ 5.1) demonstrate the existence of such an OOD prototype that exhibits strong discriminative properties. The framework itself utilizes gradient information with respect to this OOD prototype to capture distributional shifts that may not be apparent in feature representations alone.
The method comprises two primary components, illustrated in Figure 1: (1) A gradient computation framework that quantifies sample responses to an OOD prototype (§ 4.1), and (2) A nearest-neighbor scoring mechanism operating in gradient space (§ 4.2).
@@ 4.1 Gradients Computation
Inspired by the neural collapse property (Papyan et al., 2020) and prototype-based recognition (Shu et al., 2019), we construct a distance-based classification framework that incorporates both class and OOD prototypes. The key idea is to transform distances to prototypes into logits, with the crucial addition of an OOD prototype that serves as a reference point for out-of-distribution detection.
@@
For a given feature vector $h$ and a class $y\in{1,..,C,C+1}$, the cross-entropy loss associated with the NCP output $[p_{i}(h)]_{i=1}^{C+1}$ from equation 3 is given by:
@@
The key insight of our method lies in analyzing the gradient of the loss $H(h,y)$ for some (any) iid class $y$ with respect to the OOD prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$. Intuitively, this quantity represents the update vector for the OOD prototype assuming that the feature $h$ corresponds to an iid sample. This gradient can be expressed in closed form as follows:
@@
This gradient expression reveals two important properties (see the Appendix for derivation details): 1) The gradient norm equals the softmax probability $p_{\text{ood}}(h)$ of the OOD class, providing a direct measure of out-of-distribution likelihood. 2) The gradient direction, given by the unit vector $(h-\mathbf{p}^{\text{pen}}_{\text{ood}})/\|h-\mathbf{p}^{\text{pen}}_{\text{ood}}\|_{2}$, captures the geometric relationship between the sample and the OOD prototype.
@@
Figure 2: t-SNE Visualizations of ID and OOD Samples. (a) Distribution within the feature representation space including the OOD prototype using synthetic data. (b) Gradient distributions from a CIFAR-10 pre-trained ResNet-18 model, differentiating ID and OOD samples.
To analyze the potential for enhanced separability, fig. 2 visualizes (a) the penultimate feature space and (b) the gradient space using t-SNE. Our intuition is that the gradient $\nabla H(h)$ will yield a more distinct separation between ID and OOD samples compared to the feature space. As seen in fig. 2 (a), ID samples form class-specific clusters, while OOD samples are scatttered in the space. However, the t-SNE plot of the gradient space in fig. 2 (b) reveals a different spatial arrangement, suggesting that the gradient transformation highlights discriminative characteristics for OOD detection, which GROOD leverages through distance computations in this space (eq. 6).
@@ 4.2.1 Distance-Based Scoring
The preceding section outlined the core methodology of GROOD, detailing how gradients with respect to an OOD prototype are computed and subsequently used within a nearest-neighbor scoring mechanism to distinguish ID and OOD samples. A critical element underpinning the effectiveness of this methodology is the choice and construction of the OOD prototype, $\mathbf{p}^{\text{pen}}_{\text{ood}}$. The nature and location of this prototype in the feature space directly influence the direction and magnitude of the computed gradients, thereby impacting the separability of ID and OOD samples in the gradient space. Therefore, the following section delves into the specific strategies we employ for Prototype Computation (§ 5), exploring various approaches to define both the class prototypes and, crucially, the OOD prototype. These approaches are designed to yield a $\mathbf{p}^{\text{pen}}_{\text{ood}}$ that optimizes the discriminative power of the gradient-based OOD score introduced in our methodology.
@@ 5.1 Class and OOD Prototypes
Table 1: Oracle experiment results compared to SOTA excluding GROOD from Table 2. AUROC (%) on far-OOD and near-OOD detection using 100 prototype samples. Results are averaged over different checkpoints; standard deviations in parentheses.
@@
Local Oracle (Idealized): For each ID-OOD test pair, we used a small sample (100) from the test OOD data to build a specific OOD prototype. We then tested the remaining OOD data. This setup essentially asks: “If we had perfect knowledge of a small subset of the specific OOD data we’d encounter, how well could our method perform?”. The remarkable performance achieved in this setting (over $95\%$ AUROC on far-OOD detection, as shown in table 1) highlights the inherent potential of an OOD prototype tailored to the specific distributional shift.
@@ 5.2 Practical OOD Prototype Construction
The oracle experiments presented in § 5.1 demonstrated the significant potential of employing an OOD prototype to distinguish between ID and OOD data, achieving high performance when even approximate knowledge of the OOD distribution was available. This motivates our goal: to effectively approximate such an optimal OOD prototype, $\mathbf{p}^{\text{pen}}_{\text{ood}}$, without requiring access to the specific test OOD distribution, which is unavailable in practical scenarios.
While real-world OOD samples exhibit considerable diversity, representing them with a single prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$ proves effective within our gradient-aware framework. Our approach does not aim to represent all OOD samples geometrically, but rather uses the prototype as a fixed reference point. The core OOD score relies on the sensitivity of the NCP loss to this prototype, measured via the gradient $\nabla_{\mathbf{p}^{\text{pen}}_{\text{ood}}}H(h)$ (§ 4, eq. 5). We hypothesize that OOD samples, inherently deviating from learned ID manifolds, exhibit distinct gradient sensitivity patterns (both magnitude and direction) relative to this OOD reference point, allowing separation from more stable ID samples.
@@
where $\lambda=0.5$ positions the synthetic samples near decision boundaries. This approach leverages our observation that early layer representations are more sensitive to perturbations, making them ideal for generating OOD-like features.
@@
When available, we can utilize a small auxiliary OOD validation set as $X_{ood}$ to construct the prototype following eq. 8 using 100 OOD validation samples. Importantly, we ensure these samples have no category overlap with the test set. Our method shows remarkable stability to the specific choice of validation samples, with a maximum AUROC standard deviation of only 0.5% across five different random selections.
@@
A comparison of alternative $X_{OOD}$ generation methods is presented in the appendix. The results reported in the remainder of this paper utilize the OOD prototype derived from synthetic data generated from ID samples.
@@ 6 Experiments
For a comprehensive evaluation of GROOD’s performance, we adhere to the OpenOOD v1.5 criteria Zhang et al. (2023a); Yang et al. (2022). Results are aggregated in table 2 including the performance of the nearest class prototype used instead of the classification head. For robustness, each evaluation metric except for ImageNet-1k is derived from three runs with unique initialization seeds. In the case of ImageNet-1k, we report results based on a single seed run provided by torchvision maintainers & contributors (2016).
@@
We evaluate performance using the Area Under the Receiver Operating Characteristic curve (AUROC), where higher values are better. Our benchmarking strategy follows the OpenOOD framework Zhang et al. (2023a); Yang et al. (2022), involving four core ID datasets (CIFAR-10, CIFAR-100, ImageNet-200, ImageNet-1k) and examining both near and far-OOD scenarios. For CIFAR-10/100 (50k train/10k test images each), near-OOD datasets are CIFAR-100/TinyImageNet, and far-OOD are MNIST, SVHN, Textures, and Places365. For ImageNet-200 (200 classes, 64x64 resolution), near-OOD datasets are SSB-hard/NINCO, and far-OOD are iNaturalist, Textures, and OpenImage-O; ImageNet-1k shares these OOD datasets. Regarding configuration, we deploy ResNet-18 for CIFAR-10/100 and ImageNet-200 using pre-trained checkpoints from OpenOOD for consistency, testing with three distinct seeds for robustness. For ImageNet-1k, we apply pre-trained torchvision models (ResNet-50, ViT-B-16, Swin-T) to explore GROOD’s effectiveness in a broader context than OpenOOD v1 Yang et al. (2022). To allow for reproducibility and facilitate further research, the complete code, including training and evaluation scripts, is available at: https://anonymous.4open.science/r/grood.
Table 2: Main results from OpenOOD v1.5 on standard OOD detection (AUROC). GROOD using synthetic OOD data(§ 5.2) shows superior results compared to existing baselines.
@@
The main OOD detection results, presented as AUROC (%) in table 2, demonstrate the effectiveness of GROOD compared to existing state-of-the-art methods across standard benchmarks. GROOD achieves top performance in several categories, notably securing the best results for Near-OOD detection on CIFAR-10 (91.16%), ImageNet-200 (83.4%), and ImageNet-1K (78.91%), as well as for Far-OOD detection on CIFAR-100 (84.44%). In other cases, such as Far-OOD detection on CIFAR-10 (93.8%), ImageNet-200 (92.19%), and ImageNet-1k (94.57%), GROOD ranks as the second-best performing method, closely trailing the leader. Overall, the results indicate that GROOD consistently achieves competitive or superior performance across diverse datasets and network architectures (as explored further in table 3), highlighting its robustness and validating the efficacy of leveraging gradient-space geometry with respect to an OOD prototype for enhanced OOD detection.
@@
Further demonstrating the robustness and generalization capabilities of our approach, table 3 presents the OOD detection performance on ImageNet-1K using Transformer-based architectures, ViT-B-16 (Dosovitskiy et al., 2020) and Swin-T (Liu et al., 2021b). GROOD maintains strong performance, achieving the top AUROC scores for both Near-OOD and Far-OOD detection on ViT-B-16 (Dosovitskiy et al., 2020), and the best Near-OOD score on Swin-T (Liu et al., 2021b) while being highly competitive for Far-OOD. This contrasts significantly with several other methods, such as GradNorm and ASH, whose performance severely degrades on these Transformer architectures compared to their ResNet-based results. This suggests that GROOD’s mechanism, relying on gradient sensitivity relative to class and OOD prototypes, generalizes more effectively across fundamentally different architectural paradigms than methods potentially more sensitive to specific CNN feature properties.

### v3

@@ 1 Introduction
To address these challenges, we propose a novel post-hoc method called GRadient-aware Out-Of-Distribution detection (GROOD) that relies on feature and gradient spaces for improved OOD discrimination addressing three persistent challenges. First, it improves the detection of near-OOD samples those that are semantically close to the in-distribution by leveraging gradient vectors with respect to an artificial OOD prototype, which provide a more discriminative signal than feature or logit-based scores. Second, GROOD generalizes robustly across diverse architectures, including ResNets and Vision Transformers, where many existing methods suffer performance degradation. Third, it significantly reduces hyper-parameter sensitivity and exhibits stable AUROC performance across training epochs, alleviating the common issue of checkpoint instability. These properties make GROOD a practical and reliable post-hoc framework for real-world deployment.
@@
Our method is inspired by two complementary observations. First, the neural collapse (NC) property (Papyan et al., 2020) suggests that, at the end of neural network training, within-class variability tends to zero for sample representations in the feature space. This motivates the use of nearest class prototype (NCP) classification, which relies on distances to class prototypes, defined as the means of the samples of each class in the feature space. To further enhance the discriminative power of this approach for out-of-distribution detection, we extend the logits to also incorporate distances to an additional OOD prototype, alongside the distances to the class prototypes (§˜4.1).
@@
Second, we observe that OOD samples tend to exhibit a more dispersed distribution in the feature space compared to ID samples. Capitalizing on this characteristic, we introduce an artificial OOD prototype, strategically positioned to be distinct from the ID class prototypes. By then examining how sample representations respond to this OOD prototype, specifically by analyzing gradients of the NCP loss with respect to it, we can gain a more nuanced understanding of the differences between ID and OOD samples, leading to more effective discrimination in the gradient space.
@@
Our approach differs from traditional post-hoc methods by computing gradients with respect to an artificial OOD prototype rather than the network’s parameters. The magnitude of these gradients serves as a key indicator: for ID data, the OOD prototype has a relatively small influence on the confidence of the prediction in the feature space, reflecting stable classification. In contrast, for OOD data, the OOD prototype has a more substantial influence on the confidence of the prediction; meaning that a smaller shift in the OOD prototype’s representation is sufficient to cause a larger change in the classification confidence. We conduct an extensive empirical study following the recent methodology introduced in the OpenOOD Benchmark (Zhang et al., 2023a), but we also evaluate our method on other recent architectures.
@@
While NC is often seen as a limitation for OOD detection, as tightly clustered ID features can cause overconfident misclassifications, GROOD turns this into a strength. Exploiting the geometric regularity of class prototypes, it detects OOD samples not via raw confidence but through their abnormal gradient sensitivity to a fixed synthetic OOD prototype. This prototype does not aim to represent the full diversity of OOD data, but serves as a consistent reference point, inducing discriminative gradient responses that separate ID and OOD samples while avoiding the typical overconfidence failure mode.
@@ 3.2 Problem Setting: Out-of-Distribution Detection
$G_{\tau}(x)=\begin{cases}\text{\acs{ID}},&\text{if }S(x)\leq\tau,\\ \text{\acs{OOD}},&\text{if }S(x)>\tau,\end{cases}$
@@ 4 GROOD Methodology
Figure 1: Initially, we build ID class prototypes as the means of the activations of ID data along with an OOD prototype capturing OOD characteristics (§˜5). Subsequently, gradients of the softmax loss built upon the NCPs distance as logits are computed w.r.t. OOD prototype (§˜4.1). Finally, the OOD score is determined using nearest neighbor distance in the gradient space (§˜4.2).
@@
In this section, we introduce our proposed method GRadient-aware Out-Of-Distribution detection (GROOD), a novel framework for distinguishing between ID and OOD samples. To illustrate the core mechanism, We assume the existence of an OOD prototype (see §˜5.1), and use it to compute gradients and define our OOD score.
The method comprises two primary components, illustrated in Figure˜1: (1) A gradient computation framework that quantifies sample responses to an OOD prototype (§˜4.1), and (2) A nearest-neighbor scoring mechanism operating in gradient space (§˜4.2).
@@ 4.1 Gradients Computation
Guided by the observations in §˜1, we build a distance-based classification framework that integrates both class prototypes and an artificial OOD prototype. Under the NC property (Papyan et al., 2020), ID features concentrate around their class prototypes, making distances a natural choice for logits. In contrast, the OOD prototype is positioned away from these clusters and used as a reference point, allowing us to capture the difference between ID and OOD through their gradients.
@@
For a given feature vector $h$ and a class $y\in{1,..,C,C+1}$, the cross-entropy loss associated with the NCP output $p_{i}(h)]_{i=1}^{C+1}$ from equation 3 is given by:
@@
The key insight of our method lies in analyzing the gradient of the loss $H(h,y)$ for some (any) iid class $y$ with respect to the OOD prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$. Intuitively, this quantity represents the update vector for the OOD prototype assuming that the feature $h$ corresponds to an iid sample. As derived in §˜A.6, this gradient can be expressed in closed form as follows:
@@
These computations are grounded in two core observations. The first is inspired by the Neural Collapse phenomenon (Papyan et al., 2020), which shows that well-trained networks often align penultimate-layer features with their corresponding class prototypes. The second is that OOD samples tend to lie outside these tight clusters, often in low-density or dispersed regions. To leverage these properties, we adopt the $\ell_{2}$ distance as our metric. This choice is not arbitrary: it ensures that the gradient simplifies to the closed-form in equation 5, where the gradient’s norm directly equals the OOD softmax probability $p_{\text{ood}}(h)$, and its direction $(h-\mathbf{p}^{\text{pen}}_{\text{ood}})/\|h-\mathbf{p}^{\text{pen}}_{\text{ood}}\|_{2}$ encodes the feature’s geometric deviation from the prototype. Such a direct coupling between magnitude, direction, and OOD likelihood is essential to GROOD’s mechanism and would not arise with alternative metrics like cosine similarity.
@@
Figure 2: t-SNE plots on CIFAR-10 with ResNet-18. (a) Feature space: ID clusters vs. dispersed OOD. (b) Gradient space: clearer ID/OOD separation.
To analyze the potential for enhanced separability, fig.˜2 visualizes (a) the penultimate feature space and (b) the gradient space using t-SNE. Our intuition is that the gradient $\nabla H(h)$ will yield a more distinct separation between ID and OOD samples compared to the feature space. As seen in fig.˜2 (a), ID samples form class-specific clusters, while OOD samples are scatttered in the space. However, the t-SNE plot of the gradient space in fig.˜2 (b) reveals a different spatial arrangement, suggesting that the gradient transformation highlights discriminative characteristics for OOD detection, which GROOD leverages through distance computations in this space (eq.˜6).
@@ 4.2.1 Distance-Based Scoring
The preceding section outlined the core methodology of GROOD, detailing how gradients with respect to an OOD prototype are computed and subsequently used within a nearest-neighbor scoring mechanism to distinguish ID and OOD samples. A critical element underpinning the effectiveness of this methodology is the choice and construction of the OOD prototype, $\mathbf{p}^{\text{pen}}_{\text{ood}}$. The nature and location of this prototype in the feature space directly influence the direction and magnitude of the computed gradients, thereby impacting the separability of ID and OOD samples in the gradient space. Therefore, the following section delves into the specific strategies we employ for Prototype Computation (§˜5), exploring various approaches to define both the class prototypes and, crucially, the OOD prototype. These approaches are designed to yield a $\mathbf{p}^{\text{pen}}_{\text{ood}}$ that optimizes the discriminative power of the gradient-based OOD score introduced in our methodology.
@@ 5.1 Class and OOD Prototypes
Table 1: Oracle experiment results compared to SOTA excluding GROOD from Table˜2. AUROC (%) on far-OOD and near-OOD detection using 100 prototype samples. Results are averaged over different checkpoints; standard deviations in parentheses.
@@
Local Oracle (Idealized): For each ID-OOD test pair, we used a small sample (100) from the test OOD data to build a specific OOD prototype. We then tested the remaining OOD data. This setup essentially asks: “If we had perfect knowledge of a small subset of the specific OOD data we’d encounter, how well could our method perform?”. The remarkable performance achieved in this setting (over $95\%$ AUROC on far-OOD detection, as shown in table˜1) highlights the inherent potential of an OOD prototype tailored to the specific distributional shift.
@@ 5.2 Practical OOD Prototype Construction
The oracle experiments presented in §˜5.1 demonstrated the significant potential of employing an OOD prototype to distinguish between ID and OOD data, achieving high performance when even approximate knowledge of the OOD distribution was available. This motivates our goal: to effectively approximate such an optimal OOD prototype, $\mathbf{p}^{\text{pen}}_{\text{ood}}$, without requiring access to the specific test OOD distribution, which is unavailable in practical scenarios.
While real-world OOD samples exhibit considerable diversity, representing them with a single prototype $\mathbf{p}^{\text{pen}}_{\text{ood}}$ proves effective within our gradient-aware framework. Our approach does not aim to represent all OOD samples geometrically, but rather uses the prototype as a fixed reference point. The core OOD score relies on the sensitivity of the NCP loss to this prototype, measured via the gradient $\nabla_{\mathbf{p}^{\text{pen}}_{\text{ood}}}H(h)$ (§˜4, eq.˜5). We hypothesize that OOD samples, inherently deviating from learned ID manifolds, exhibit distinct gradient sensitivity patterns (both magnitude and direction) relative to this OOD reference point, allowing separation from more stable ID samples.
@@
where $\lambda=0.5$ positions the synthetic samples near decision boundaries. This approach leverages our observation that early layer representations are more sensitive to perturbations, making them ideal for generating OOD-like features. Although effective, mixup-generated OOD samples interpolate between ID classes and may not fully capture OOD diversity.
@@
When available, we can utilize a small auxiliary OOD validation set as $X_{ood}$ to construct the prototype following eq.˜8 using 100 OOD validation samples. Importantly, we ensure these samples have no category overlap with the test set. Our method shows remarkable stability to the specific choice of validation samples, with a maximum AUROC standard deviation of only 0.5% across five different random selections.
@@
A comparison of alternative $X_{OOD}$ generation methods is presented in the §˜A.2. The results reported in the remainder of this paper utilize the OOD prototype derived from synthetic data generated from ID samples.
@@ 6 Experiments
For a comprehensive evaluation of GROOD’s performance, we adhere to the OpenOOD v1.5 criteria Zhang et al. (2023a); Yang et al. (2022). Results are aggregated in table˜2 including the performance of the nearest class prototype used instead of the classification head. For robustness, each evaluation metric except for ImageNet-1k is derived from three runs with unique initialization seeds. In the case of ImageNet-1k, we report results based on a single seed run provided by torchvision maintainers & contributors (2016).
@@
We evaluate performance using the Area Under the Receiver Operating Characteristic curve (AUROC), where higher values are better. Our benchmarking strategy follows the OpenOOD framework Zhang et al. (2023a); Yang et al. (2022), involving four core ID datasets (CIFAR-10, CIFAR-100, ImageNet-200, ImageNet-1k) and examining both near and far-OOD scenarios. For CIFAR-10/100 (50k train/10k test images each), near-OOD datasets are CIFAR-100/TinyImageNet, and far-OOD are MNIST, SVHN, Textures, and Places365. For ImageNet-200 (200 classes, 64x64 resolution), near-OOD datasets are SSB-hard/NINCO, and far-OOD are iNaturalist, Textures, and OpenImage-O; ImageNet-1k shares these OOD datasets. Regarding configuration, we deploy ResNet-18 for CIFAR-10/100 and ImageNet-200 using pre-trained checkpoints from OpenOOD for consistency, testing with three distinct seeds for robustness. For ImageNet-1k, we apply pre-trained torchvision models (ResNet-50, ViT-B-16, Swin-T) to explore GROOD’s effectiveness in a broader context than OpenOOD v1 Yang et al. (2022). To allow for reproducibility and facilitate further research, the complete code, including training and evaluation scripts, is available at: https://github.com/mostafaelaraby/Gradient-Aware-OOD-Detection.
For CIFAR-10/100 and ImageNet-200, we train ResNet-18 models for 100 epochs using SGD with momentum 0.9, weight decay 5e-4, cosine learning rate decay (starting from 0.1), and batch sizes of 128 (CIFAR) and 256 (ImageNet-200). For ImageNet-1K, we use pretrained models from torchvision (ResNet-50, ViT, Swin). When fine-tuning is required, we follow the OpenOOD v1.5 protocol (Zhang et al., 2023a) with 30 epochs, learning rate 0.001, and batch size 256.
Table 2: Main results from OpenOOD v1.5 on standard OOD detection (AUROC). GROOD using synthetic OOD data (§˜5.2) shows superior results compared to existing baselines.
@@
GROOD shows strong performance across datasets, but performance varies depending on the trade-off between Near and Far-OOD detection. On CIFAR-100, GROOD achieves state-of-the-art Far-OOD AUROC (84.44%) among post-hoc methods and competitive Near-OOD performance (78.9%). While some methods like VIM (Wang et al., 2022) (81.70%) report higher Near-OOD scores, they trade off Far-OOD robustness.
On ImageNet-1k, GROOD achieves a top Far-OOD score (94.8%) but a lower Near-OOD score (78.91%), whereas CombOOD (Rajasekaran et al., 2024) yields 95.22% Near-OOD and 90.24% Far-OOD.
These results reflect a key characteristic of GROOD: its effectiveness depends on the structure of the ID feature space. As GROOD relies on geometric separation of class prototypes (inspired by Neural Collapse (Papyan et al., 2020)), its performance can degrade when ID representations are less well-clustered. Additionally, the choice of OOD prototype impacts this trade-off. For example, using an “ID-corrupted val” prototype improves Near-OOD AUROC to 80.27% (CIFAR-100) and 83.5% (ImageNet-1k), while maintaining strong Far-OOD scores.
Importantly, many top-performing methods on the OpenOOD leaderboard require access to OOD data (e.g., OE, CIDER). GROOD remains fully post-hoc and training-free, making it more practical for deployment across varied scenarios.
@@
Further demonstrating the robustness and generalization capabilities of our approach, table˜3 presents the OOD detection performance on ImageNet-1K using Transformer-based architectures, ViT-B-16 (Dosovitskiy et al., 2020) and Swin-T (Liu et al., 2021b). GROOD maintains strong performance, achieving the top AUROC scores for both Near-OOD and Far-OOD detection on ViT-B-16 (Dosovitskiy et al., 2020), and the best Near-OOD score on Swin-T (Liu et al., 2021b) while being highly competitive for Far-OOD. This contrasts significantly with several other methods, such as GradNorm and ASH, whose performance severely degrades on these Transformer architectures compared to their ResNet-based results. This suggests that GROOD’s mechanism, relying on gradient sensitivity relative to class and OOD prototypes, generalizes more effectively across fundamentally different architectural paradigms than methods potentially more sensitive to specific CNN feature properties.